# Law of Vision Representation in MLLMs

### v2

@@ Abstract
We introduce the “Law of Vision Representation” in multimodal large language models (MLLMs), revealing a strong correlation among cross-modal alignment, vision representation correspondence, and overall model performance. We quantify the these factors using the cross-modal Alignment and Correspondence scores. Extensive experiments across fifteen distinct vision representation settings and evaluations on eight benchmarks show that A and C scores correlate with performance following a quadratic relationship. By leveraging this relationship, we can identify and train the optimal vision representation for an MLLM, achieving a 99.7% reduction in computational cost without the need for repeated finetuning of the language model. The code is available at https://github.com/bronyayang/Law_of_Vision_Representation_in_MLLMs.
@@ 1 Introduction
Current multimodal large language models (MLLMs) (Chen et al., 2024a; Liu et al., 2024e; d) have achieved remarkable advancements by integrating pretrained vision encoders with powerful language models (Touvron et al., 2023; Zheng et al., 2023). Among the core components of a general MLLM, vision representation plays a critical role. Many researchers have utilized CLIP (Radford et al., 2021) or SigLIP (Zhai et al., 2023b; Tschannen et al., 2025) as the primary image feature encoder, but their limitations are becoming increasingly noticed (Tong et al., 2024b; Geng et al., 2023; Yao et al., 2021). As a result, alternative vision representations (Tang et al., 2025) and the combination of multiple vision encoders are being actively explored (Tong et al., 2024a; Lin et al., 2023).
@@
To address this gap in understanding what makes a vision representation optimal for MLLMs, we propose the Law of Vision Representation in MLLMs. It aims to explain the key factors of vision representation that impact MLLM benchmarks performance. Our findings reveal that cross-modal Alignment (A) and Correspondence (C) of the vision representation are strongly correlated with model performance.; specifically, higher A and C lead to improved performance. To quantify this relationship, we define A and C scores that measure cross-modal alignment and correspondence in vision representation. The A and C score as well as model performance exhibit a quadratic relationship, with a coefficient of determination of 94.06%.
@@ 3.3 Empirical Justification
    A $\textsc{A Score}(I,y)=\frac{1}{T}\sum_{t=1}^{T}\log P\big{(}y_{t}\mid f(I),y_{<t}\big{)}$
@@
    C $\textsc{C Score}=\frac{1}{m}\sum^{m}_{i=0}\mathbbm{1}_{\left\|p^{pred}_{i}-p^{GT}_{i}\right\|_{2}<T}$ (2)
@@
AC $\textsc{AC Score}=\sum^{2}_{\alpha=0}\sum^{2-\alpha}_{\beta=0}w_{\alpha\beta}A^{\alpha}C^{\beta}$ (3)
@@ Results.
[…]
tableVision representations explored.
@@ 5.1 Experiment Settings
For our MLLM pipeline, we deploy a LLaMA-based LLM, specifically Vicuna-7B 1.5 (Zheng et al., 2023), and utilize a widely used 2-layer GeLU-MLP connector as the projector. For vision representation, we explore a variety of encoder types and sizes, input resolutions, training paradigms, and feature combinations, as detailed in Table 4.
@@ 5.4 Additional Results on the Law of Vision Representation
In Section 3, we demonstrate the strong correlation between the AC score and MLLM performance by analyzing the coefficient of determination ($R^{2}$) obtained from fitting a quadratic regression model. In this section, we further ablate the experiments by adding baselines, fitting model performance with random scores, A scores, and C scores separately. Additionally, we explored the relationship between the A score and C score by fitting a linear regression model. We avoid higher-degree transformations to prevent overfitting, which could obscure the true relationship between A and C scores.
@@ A.1 Theoretical Justification of Vision Representation with High Cross-modal Alignment
where $\epsilon$ is a small constant. Given this condition, we can show that the output of the MLLM with multimodal embeddings $[E_{c}^{image},E_{1},E_{2},\dots,E_{n}]$ is close to the output with text-only embeddings $[E_{c}^{text},E_{1},E_{2},\dots,E_{n}]$.
@@ A.2 Theoretical Justification of Vision Representation with Accurate Correspondence
Consider an input $[E_{0}^{\text{image}},E_{1}^{\text{image}},E_{2},\dots,E_{n}]$ to the transformer, where the image embeddings $E_{0}^{\text{image}}$ and $E_{1}^{\text{image}}$ are derived from different patch of a high correspondence vision representation. By definition, the dot product $E_{0}^{\text{image}}\cdot E_{1}^{\text{image}}$ is large if the two corresponding original image patches share related information.
@@
A.3 All Settings Benchmark Performance
@@ A.4 All Settings AC Scores
A.5 More Visualization of Correspondence
@@ A.6 Pseudo Code
Input: $k$ A and C score pairs from models $ACs$; $past\_sampled$ models; current sampling $level$ ($1$ to $k^{\prime}$, increments when regions are exhausted as each region is sampled only once)
Output: Sampled $model$ to train next
@@
       $regions[region\_key].append((model,A,C))$;
Remove models in $past\_sampled$ from $regions$;
@@ A.7 Limitation of AC Policy
Figure 9: AC policy.
A.8 AC Policy Recall@1
In Section 5.5, we used Recall@3 as a metric for AC policy effectiveness to capture potential performance fluctuations during MLLM training and evaluation. Here, we also report Recall@1 for predicting the optimal vision representation. As shown in Figure 10, AC policy consistently outperforms random testing even under this stricter metric.

### v3

@@ Abstract
We introduce the “Law of Vision Representation” in multimodal large language models (MLLMs), revealing a strong correlation among cross-modal alignment, vision representation correspondence, and overall model performance. We quantify these factors using the cross-modal Alignment and Correspondence scores. Extensive experiments across fifteen distinct vision representation settings and evaluations on eight benchmarks show that A and C scores correlate with performance following a quadratic relationship. By leveraging this relationship, we can identify and train the optimal vision representation for an MLLM, achieving a 99.7% reduction in computational cost without the need for repeated finetuning of the language model. The code is available at https://github.com/bronyayang/Law_of_Vision_Representation_in_MLLMs.
@@ 1 Introduction
Current multimodal large language models (MLLMs) (Chen et al., 2024a; Liu et al., 2024e; d) have achieved remarkable advancements by integrating pretrained vision encoders with powerful language models (Touvron et al., 2023; Zheng et al., 2023). Among the core components of a general MLLM, vision representation plays a critical role. Many researchers have utilized CLIP (Radford et al., 2021) or SigLIP (Zhai et al., 2023b; Tschannen et al., 2025) as the primary image feature encoder, but their limitations are becoming increasingly noticeable (Tong et al., 2024b; Geng et al., 2023; Yao et al., 2021). As a result, alternative vision representations (Tang et al., 2025) and the combination of multiple vision encoders are being actively explored (Tong et al., 2024a; Lin et al., 2023).
@@
To address this gap in understanding what makes a vision representation optimal for MLLMs, we propose the Law of Vision Representation in MLLMs. It aims to explain the key factors of vision representation that impact MLLM benchmark performance. Our findings reveal that cross-modal Alignment (A) and Correspondence (C) of the vision representation are strongly correlated with model performance.; specifically, higher A and C lead to improved performance. To quantify this relationship, we define A and C scores that measures cross-modal alignment and correspondence in vision representation. The A and C scores as well as model performance exhibit a quadratic relationship, with a coefficient of determination of 94.06%.
@@ 3.3 Empirical Justification
    $\textsc{A Score}(I,y)=\frac{1}{T}\sum_{t=1}^{T}\log P\big(y_{t}\mid f(I),y_{<t}\big)$
@@
    $\textsc{C Score}=\frac{1}{m}\sum^{m}_{i=0}\mathbbm{1}_{\left\|p^{pred}_{i}-p^{GT}_{i}\right\|_{2}<T}$ (2)
@@
$\textsc{AC Score}=\sum^{2}_{\alpha=0}\sum^{2-\alpha}_{\beta=0}w_{\alpha\beta}A^{\alpha}C^{\beta}$ (3)
@@ Results.
Table 1: Vision representations explored.
@@ 5.1 Experiment Settings
For our MLLM pipeline, we deploy LLaMA-based LLM, specifically Vicuna-7B 1.5 (Zheng et al., 2023) as well as Qwen-14B 2.5 Instruct (Qwen et al., 2025), and utilize a widely used 2-layer GeLU-MLP connector as the projector. For vision representation, we explore a variety of encoder types and sizes, input resolutions, training paradigms, and feature combinations, as detailed in Table 4.
@@ 5.4 Additional Results on the Law of Vision Representation
In Section 3, we demonstrate the strong correlation between the AC score and MLLM performance by analyzing the coefficient of determination ($R^{2}$) obtained from fitting a quadratic regression model. In this section, we further ablate the experiments by adding baselines, fitting model performance with random scores, A scores, and C scores separately. Additionally, we explored the relationship between the A score and C score by fitting a linear regression model. We provide more insights about the AC relationship in appendix A.3. Besides, we avoid higher-degree transformations to prevent overfitting, which could obscure the true relationship between A and C scores.
@@
LLM Backbone $R^{2}$ (Vision) $R^{2}$ (OCR)
Vicuna 1.5 7B 94.06% 83.85%
Qwen 2.5 14B 91.50% 82.27%
Table 3: Averaged law fitting $R^{2}$ comparison for different types and sizes of LLM backbones.
Here, we additionally provide MLLM with Qwen 2.5 14B Instruct as LLM backbone fitting results, as shown in Table 3. Specifically, we use the exact same settings (data, vision representations, hyperparameters), changing only the LLM backbone to Qwen 2.5 14B. We again evaluate the resulting MLLMs on 8 benchmarks, recompute the A and C scores, reporting the fitting $R^{2}$. The averaged fitting $R^{2}$ are above 90% for vision based benchmarks and above 80% for OCR based benchmarks, comparable to Vicuna 1.5 7B, showing that our conclusion applies to different types and sizes of LLM backbones.
@@ 5.5 Additional Results on the AC Policy
Obtaining the AC policy results in significant computational cost reduction. After fitting the AC function, testing a new vision encoder only requires MLLM Stage 1 training (i.e. training a small projector), while traditional full MLLM post-training requires training both the projector and LLM. Concretely, we quantify computational cost using the number of trainable parameters, which is commonly correlated with both memory and compute. A decoder-only MLLM’s Stage 1 trains a small two-layer MLP, while Stage 2 involves fine-tuning a large language model with 7B parameters. This leads to a parameter ratio of about 0.003, yielding a 99.7% reduction.
@@ 6 Limitation
7 Conclusion
In this work, we present the Law of Vision Representation in MLLMs, uncovering a strong and quantifiable relationship between cross-modal alignment, feature correspondence, and overall model performance. Our proposed A and C scores capture this relationship with high fidelity, enabling a principled way to reason about vision representation quality. Through extensive evaluation, we demonstrate that model performance exhibits a clear quadratic correlation with AC scores, providing not only an interpretable diagnostic but also a predictive signal for selecting optimal visual features.
By leveraging this insight, we introduce the AC policy, a cost-efficient strategy to identify high-performing vision representations without repeatedly finetuning the language model. This yields up to 99.7% reduction in compute cost, transforming what was previously an expensive and combinatorially large search problem into a tractable one. Our findings offer a new perspective on vision-language alignment and open the door to more scalable MLLM development.
@@ A.1 Theoretical Justification of Vision Representation with High Cross-modal Alignment
where $\epsilon$ is a small constant. Given this condition, we can show that the output of the MLLM with multimodal embeddings $E_{c}^{image},E_{1},E_{2},\dots,E_{n}]$ is close to the output with text-only embeddings $E_{c}^{text},E_{1},E_{2},\dots,E_{n}]$.
@@ A.2 Theoretical Justification of Vision Representation with Accurate Correspondence
Consider an input $E_{0}^{\text{image}},E_{1}^{\text{image}},E_{2},\dots,E_{n}]$ to the transformer, where the image embeddings $E_{0}^{\text{image}}$ and $E_{1}^{\text{image}}$ are derived from different patch of a high correspondence vision representation. By definition, the dot product $E_{0}^{\text{image}}\cdot E_{1}^{\text{image}}$ is large if the two corresponding original image patches share related information.
@@
A.3 Analysis of Quadratic Relationship between A and C
We first provide an intuitive explanation of why we choose to capture a quadratic relationship between cross-modal alignment and correspondence. Then, we analysis the coefficent of the fitted function to reveal a more in-depth relationship between these two factors.
Here’s our insight:
  • Good semantic correspondence means the vision part of MLLM can really tell what things are in a picture, even if they look a bit different. Like knowing all dogs are dogs, no matter their breed. But this alone isn’t enough for the MLLM to ”talk” about them. Those visual details need to be understood by the LLM.
  • Good language alignment means the visual information is perfectly understood by the language model. But if the initial visual information isn’t very good at telling different things apart (poor semantic correspondence), then aligning it perfectly won’t help much when the MLLM sees something new or slightly different.
So, for MLLMs to work best, both semantic correspondence and language alignment need to work together. More importantly, they interplay with each other more than just adding up:
If the visual features are already well-aligned with language (like in models such as CLIP), it actually helps the MLLM understand the semantic details better. That’s because if you can map ”dog” to the word ”dog” consistently, it reinforces the concept of ”dog” in the visual world.
But why is this interplay represented in a quadratic form (like a hill)? Think of it in this way:
If language alignment is very weak, even if you make semantic correspondence a little better, the MLLM’s performance won’t jump much. It’s like trying to have a good conversation when you don’t speak the same language very well. Take a model like DINOv2. It’s great at seeing things, but it might not be as good at linking those visual details to language as, say, CLIP. In this case, making the visual part even better won’t help MLLM performance much because the language alignment is the bottleneck.
As language alignment improves, the MLLM can then really benefit from better semantic correspondence. It’s like they’re working together to reach that ”sweet spot” where the performance really shines. Both getting better at the same time, or one unlocking the potential of the other, leads to this curved, quadratic rise in performance.
In short, quadratic models include $A^{2}$ and $C^{2}$ and interaction term $\alpha_{AC}$ which includes exactly the relationship we want to capture. In our paper, we test linear, second-degree polynomial relationships. Our result shows that the second-degree relationship fitted the best among others while still maintaining simplicity. More complex functional forms would obscure the clear individual contribution of A and C, as well as interaction between alignment and correspondence.
Next, We discuss the interation alignment and correspondence from two aspects: the ideal relationship between A and C suggested by the law, and the actual relationship we observe from existing encoders. To provide the ideal relationship, we extract the cross-term coefficient ($\alpha_{AC}$) of the fitted functions as shown in Table 4.
Benchmark $\alpha_{AC}$
MMBench +1.1301
MME +1.5819
MMMU -0.4872
OKVQA -0.0364
TextVQA +4.9804
VizWiz +2.9964
ScienceQA +4.4644
SeedBench +0.5772
Table 4: Interaction term $\alpha_{AC}$ of fitted function for all eight benchmarks.
We observe two kind of interactions:
Synergy ($\alpha_{AC}\gg 0$): MMBench, MME, TextVQA, VizWiz, ScienceQA, and SeedBench all have $\alpha_{AC}>0$. For these six tasks, performance improves more than additively when both $A$ and $C$ increase.
Mild trade-off ($\alpha_{AC}<0$): MMMU and OKVQA show $\alpha_{AC}<0$, indicating a slight penalty if $A$ and $C$ are both pushed upward (i.e., diminishing returns).
These results demonstrate that on most benchmarks, alignment and correspondence actually reinforce each other (positive interaction). Therefore, $A$ and $C$ strengthen each other, so the ideal scenario would be increasing both to achieve the best performance.
However, from our observations of $A$ and $C$ of vision encoders, the factors often have a trade-off: Intuitively, $A$ measures how well image features align to a shared text embedding, whereas $C$ captures intra-image consistency (e.g., visual cluster compactness). Forcing very high alignment often collapses visual cluster structure (hurting correspondence). In practice, looking at the $A$ and $C$ scores in Appendix Section A.5, common visual encoders often have high $A$ but low $C$ (CLIP and SigLIP) or low $A$ but high $C$ (DINOv2 and Diffusion features).
A.4 All Settings Benchmark Performance
@@ A.5 All Settings AC Scores
A.6 More Visualization of Correspondence
@@ A.7 Pseudo Code
Input: $k$ A and C score pairs from models ACsACs; past_sampledpast\_sampled models; current sampling levellevel (11 to $k^{\prime}$, increments when regions are exhausted as each region is sampled only once)
Output: Sampled modelmodel to train next
@@
 regions[region_key].append((model,A,C))regions[region\_key].append((model,A,C));
Remove models in past_sampledpast\_sampled from regionsregions;
@@ A.8 Limitation of AC Policy
Figure 9: Limitation of AC policy.
A.9 AC Policy Recall@1
In Section 5.5, we used Recall@3 as a metric for AC policy effectiveness to capture potential performance fluctuations during MLLM training and evaluation. Our AC policy’s core aim is to narrow the search space to a small set of promising vision encoders, rather than pinpointing a single best encoder. Recall@3 directly quantifies whether the true top-performers appear within our top-3 recommendations, which:
  • Reflects practical workflows—researchers typically inspect multiple (“top 3–5”) candidates, not just one.
  • Smooths score noise—unlike Top-1 accuracy, Recall@3 tolerates minor ranking fluctuations while still ensuring coverage of strong encoders.
  • Focuses on coverage—rather than overall ranking correlation (e.g. Pearson’s $r$) or mean reciprocal rank (which over-emphasizes exact positions), Recall@3 measures actionable retrieval quality: “Is the best encoder in my shortlist?” In other words, we put coverage in higher priority, “Is the best encoder in my shortlist” is more important than is the top-3 order correct.
While we believe this makes Recall@3 the most appropriate metric for AC’s intended use-case, we also report Recall@1 for predicting the optimal vision representation. As shown in Figure 10, AC policy consistently outperforms random testing even under this stricter metric.